OpenStudy (anonymous):

The number of Math homework questions is seven more than one-half the number of History homework questions. There are 15 Math homework questions. Which of the following equations would you use to find the number of History homework questions? Note: h represents the number of questions on your History homework.

OpenStudy (hoblos):

15= h/2 + 7

OpenStudy (anonymous):

minus 7 on both sides then multiplies both sides by 2

OpenStudy (anonymous):

8 = h/2 h = 16

OpenStudy (anonymous):

.5h + 7 = 15 .5h = 8 h = 8/.5 h = 16

OpenStudy (anonymous):

16 questions

OpenStudy (hoblos):

i though he just needed the equation

OpenStudy (anonymous):

yea i just need the equation